Cabrina and David are Avon Representatives who are attending a conference in Seattle. After the conference, Cabrina drives home to Boise at an average speed of 65 miles per hour and David drives home to Portland at an average speed of 50 miles per hour. If the sum of their driving times is 11.4 hours and if the sum of the distance driven is 690 miles, determine the time each representative spent driving home.
How would I go about setting up equations for this problem? And how would I go about solving it?

Answer by checkley77(12844)

D=RT
690=65T+50(11.4-T)
690=65T+570-50T
65T-50T=690-570
15T=120
T=120/15
T=8 HOURS FOR THE 65 MPH TRIP.
11.4-8=3.4 HOURS FOR THE 50 MPH TRIP.
PROOF:
690=65*8+50*3.4
690=520+270
690=690